• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 08:03:09 on June16,2002

    Entry number 86140

    keyword=owl shift summary

    Crew: R. Michaels, O. Hansen, A. Camsonne, G. Batigne

    At start of shift we took access to power cycle the ROC14 crate. That fixed it.

    At first MCC could not deliver more than 50 uA. Later (during run 1528) they
    managed to reach 100 uA.

    Run...............Comments

    1523 & 24.......junk
    1525..............Production Run, Kinematics Q2D_d20
    1526..............Production, but the beam is very unstable and mostly absent
    1527..............brief run (2.3 min)
    1528..............Production.....Beam is 70 uA at start, ramps up to 104 uA gradually
    1529..............Production.... 1 M events.... good beam
    1530..............Efficiency study run (ps1 = ps2 = 1)

    Next, we change to Kinematics Q2D_d40 by changing R-HRS to
    P_R = 1.9983 theta_R = 48.86 (no quad cycling since we go down in P_R).

    1531...............junk (CODA problem at prestart)
    1532...............9 Foil Optics
    1533...............Production......ended with ROC14 crash.....800 K events
    1534...............Production......ROC14 again.....345 K
    1535...............Production..... 1 M
    1536...............junk, does not exist
    1537...............Production.....in progress

    The ROC14 crashes are more frequent -- partly due to increased radiation load
    (we got 100 uA), but maybe the cpu is dying (replacing it has, in the past, helped).
    If this becomes intolerable I can replace the cpu.